OpenStudy (anonymous):

Please answer ASAP will fan and medal What is the slope of the graph of 6x + 12y = 7? A. -2 B. -1/2 C. 1/2 D. 2

OpenStudy (anonymous):

y = -6/12x + 7/12 therefore the slope is -6/12 = -1/2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You re welcome

OpenStudy (anonymous):

>,<